OpenStudy (anonymous):

Factor completely: 2x^2 + 2x - 24

OpenStudy (anonymous):

\[2x^2 + 2x - 24 = 2(x^2 +x - 12)= 2( x-?)(x+?)\]

OpenStudy (anonymous):

\[2x^2 + 2x - 24 = 2(x^2 +x - 12)= 2( x-?)(x+4) \] What can you replace ? with?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

I'm not very good at factoring:/

OpenStudy (anonymous):

yes 3

OpenStudy (anonymous):

Thank you:)